Operations Report Jan 25

Crew 220 Operations Report 25-JAN-2020

Sol: 6

Name of person filing report: John Hanacek

Non-nominal systems: N/A

Generator: Not Run

Hours run: 0

From what time last night: N/A

To what time this morning: N/A

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): n/a

Diesel Reading – N/A

Station Propane Reading- 2%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – N/A

Water Meter (units)- N/A

Water (static tank) (gal) – N/A

Static to loft Pump used – N/A

Water in Green Hab (gal): N/A

Water in Science Dome (gal): N/A

Toilet Tank Emptied – no

Diemos rover used: N/A

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – N/A

Starting Hours: N/A

Beginning Charge: N/A

Ending Hours: N/A

Ending Charge: N/A

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: N/A

Starting Hours: N/A

Beginning Charge: N/A

Ending Hours: N/A

Ending Charge: N/A

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: N/A

Summary of Suit and Radios: Nominal

Summary of Hab: Structural integrity nominal. Floors becoming dirty due to mud. Will be reinforcing with cardboard

Summary of Science Dome operations: Temperatures are hovering around 7°C during the day.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None